# Goal-oriented time adaptivity for port-Hamiltonian systems

### v1

@@ 5. A dissipativity exploiting parallelizable strategy for the adjoints
$\left(\begin{smallmatrix}I-h_{1}A^{\top}&0&0&0&\cdots&0\\ 0&I-h_{2}A^{\top}&0&0&\cdots&0\\ 0&0&I-h_{3}A^{\top}&0&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&0&I-h_{M}A^{\top}&0\\ 0&0&0&0&0&I\\ \end{smallmatrix}\right)\begin{pmatrix}\tilde{\lambda}_{0}\\ \tilde{\lambda}_{1}\\ \vdots\\ \tilde{\lambda}_{M}\end{pmatrix}=\begin{pmatrix}\langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{1}\rangle_{X_{h}^{*},X_{h}}\\ \langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{2}\rangle_{X_{h}^{*},X_{h}}\\ \vdots\\ \langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{M}\rangle_{X_{h}^{*},X_{h}}\end{pmatrix}$ (5.1)
@@
Then, for all $t\in[t_{i},t_{i+1}]$
$\|\lambda(t)-\tilde{\lambda}(t)\|_{Q}\leq\|\lambda_{T}\|_{Q}.$ (5.2)
@@
Approximating the computation of the sensitivities in a (block-)diagonal manner is very closely related to the DLY (Deuflhard-Leinen-Yserentant; [6]) methodologies for elliptic problems. There, due to fast decay of the Green’s function in space, the error transport is also localized. Here, the dissipativity of the port-Hamiltonian model in time prevents the global error transport.
@@ 6. Numerical results
$J=\begin{pmatrix}0&0&1\\ 0&0&-1\\ -1&1&0\end{pmatrix},\quad R_{1}=\begin{pmatrix}1&1&0\\ 1&1&0\\ 0&0&0\end{pmatrix},\quad R_{2}=I,\quad Q=I,\quad B=\begin{pmatrix}1\\ 1\\ 0\end{pmatrix},$ (6.1)
@@ 6.1. Approximation of the sensitivities
6.2. Performance of the goal-oriented scheme
@@
First, in Figure 2, we evaluate the performance of $I_{\mathrm{loc}}$ with sensitivities computed from (4.3) with a scheme using approximated sensitivities from (5.1). For the sensitivies computed via the exact adjoint system (4.3), we observe in the bottom left plot that the violation of the energy balance is always lowest when using the goal-oriented refinement regime. We stress that odeint internally utilizes a higher number of grid points. In the upper plots of Figure 2 we observe that despite this good performance in view of the goal of refinement, the error of the scheme is relatively large in the third component towards the end of the horizon, as the refinement w.r.t. $I_{\mathrm{loc}}$ solely focuses on the error in the energy balance. In the right column of Figure 2, we plot the results using approximated sensitivities. Whereas for a smaller amount of time steps, the energy balance violation is still the lowest, we reach a point of saturation compared to odeint. Intuitively, this indicates that the dissipation is not strong enough to justify localizations to very small time steps occurring. We note that the considered matrix $A=(J-R_{1})Q$ is not Hurwitz due to the presence of two imaginary eigenvalues, that is, the system is not exponentially stable.
@@ 7. Conclusion and outlook
We proposed an adaptive and goal-oriented grid refinement approach tailored to the energy balance in port-Hamiltonian systems. To this end, we formulated the state equation and the corresponding adjoint equation as variational problems and proved a result for higher regularity of the adjoint for right-hand sides occurring in the proposed scheme. To guarantee an efficient evaluation of the error estimator, we further provided a block-diagonal and stability-exploiting approximation for the adjoint system which we motivated from a linear algebraic perspective via a block-Jacobi method. Then, we provided a numerical example highlighting that the proposed scheme leads to smaller violations of the energy balanced when compared to step-size controlled or higher-order time stepping schemes.
In the future, we plan to consider goal-oriented multirate schemes [9] for pHs and to relate the dissipativity-based arguments of Section 4 with the approach using generalized Green’s functions to capture error transport proposed in [10].

### v2

@@ 5. A dissipativity exploiting parallelizable strategy for the adjoints
$\left(\begin{smallmatrix}I-h_{1}A^{\top}&0&0&0&\cdots&0\\ 0&I-h_{2}A^{\top}&0&0&\cdots&0\\ 0&0&I-h_{3}A^{\top}&0&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&0&I-h_{M}A^{\top}&0\\ 0&0&0&0&0&I\\ \end{smallmatrix}\right)\begin{pmatrix}\tilde{\lambda}_{0}\\ \tilde{\lambda}_{1}\\ \vdots\\ \tilde{\lambda}_{M}\end{pmatrix}=\left(\begin{smallmatrix}\langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{1}\rangle_{X_{h}^{*},X_{h}}\\ \langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{2}\rangle_{X_{h}^{*},X_{h}}\\[1.99997pt] \vdots\\[6.99997pt] \langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{M}\rangle_{X_{h}^{*},X_{h}}\end{smallmatrix}\right)$ (5.1)
@@
Then, for all $t\in[t_{i},t_{i+1}]$, and denoting $\|v\|_{Q}:=\sqrt{v^{\top}Qv}$,
$\|\lambda(t)-\tilde{\lambda}(t)\|_{Q}\leq\|\lambda_{T}\|_{Q},$ (5.2)
@@
Approximating the computation of the sensitivities in a block-diagonal manner is very closely related to DLY (Deuflhard-Leinen-Yserentant; [6]) methodologies for elliptic problems. There, due to fast decay of the Green’s function in space, the error transport is also localized. Here, the dissipativity of the port-Hamiltonian model in time prevents the global error transport.
@@ 6. Numerical results
$J=\begin{pmatrix}0&0&1\\ 0&0&-1\\ -1&1&0\end{pmatrix},\ R_{1}=\begin{pmatrix}1&1&0\\ 1&1&0\\ 0&0&0\end{pmatrix},\ R_{2}=I,\ Q=I,\ B=\begin{pmatrix}1\\ 1\\ 0\end{pmatrix},$ (6.1)
@@ 6.1. Approximation of the sensitivities
6.2. Performance of the goal-oriented scheme: Toy example
@@
First, in Figure 2, we evaluate the performance of $I_{\mathrm{loc}}$ with sensitivities computed from (4.3) with a scheme using approximated sensitivities from (5.1).
@@
For the sensitivities computed via the exact adjoint system (4.3), we observe in the bottom left plot that the violation of the energy balance is always lowest when using the goal-oriented refinement regime. We stress that odeint internally utilizes a higher number of grid points. In the upper plots of Figure 2 we observe that despite this good performance in view of the goal of refinement, the error of the scheme is relatively large in the third component towards the end of the horizon, as the refinement w.r.t. $I_{\mathrm{loc}}$ solely focuses on the error in the energy balance. In the right column of Figure 2, we plot the results using approximated sensitivities. Whereas for a smaller amount of time steps, the energy balance violation is still the lowest, we reach a point of saturation compared to odeint. Intuitively, this indicates that the dissipation is not strong enough to justify localizations corresponding to the block-diagonal approximation. We note that the considered matrix $A=(J-R_{1})Q$ is not Hurwitz due to the presence of two imaginary eigenvalues, that is, the system is not exponentially stable.
@@
Last, we also provide a comparison to the implicit midpoint rule which is known to be one of the simplest symplectic and (due to the quadratic Hamiltonian) energy-preserving integrators [16, 24, 20]. We consider the case without input, that is $u\equiv 0$ and compare the performance for different dissipation matrices. In case of no dissipation, the implicit midpoint rule leads to no violation of the energy balance. In Figure 4 we depict the results. On the left we see that, for the dissipation matrix considered earlier, the implicit midpoint rule overtakes the goal-oriented adaptive implicit Euler scheme. This is due to the implicit midpoint rule enjoying a higher convergence rate by one order. On the right-hand side however, we see that for smaller dissipation, the implicit midpoint rule outperforms the goal-oriented scheme for all choices of time steps. This clearly motivates future research formulating goal-oriented adaptive midpoint schemes: The adaptivity then particularly targets time instances at which energy is dissipated (or supplied) and thus complements the symplectic properties of the standard implicit midpoint scheme.
Figure 4. Violation of energy balance and comparison to implicit midpoint rule with uniform refinement with no control input: Left: $R$ from (6.1), Right: Dissipation given by $0.6\cdot R$.
6.3. Performance of the goal-oriented scheme: Electrical circuit
In this part, we consider a circuit model in the sense of a modified nodal analysis as depicted in Figure 5.
[…]
Figure 5. Depiction of the considered electrical circuit
The variables are the node potential $e_{i}$ ($i=0,1,\dotsc,2n$), and the currents through the inductors $\jmath_{i}$ ($i=1,\dotsc,n$). The corresponding pHs formulation is given by
$E\dot{x}(t)=(J-R)x(t)+Bu(t)$
with the state variable
$x=[e_{0},e_{1},\jmath_{1},e_{2},e_{3},\jmath_{2},\dotsc,e_{2n-1},\jmath_{n},e_{2n}]^{\top}\in^{3n+1}$
the (invertible) descriptor matrix
$E=\text{blkdiag}(C,E_{0},\dotsc,E_{0})\in^{(3n+1)\times(3n+1)}\ \ \text{with}\ \ E_{0}=\text{diag}(C_{R},L,C)\in^{3\times 3},$
the structure matrix
$J=\text{blkdiag}(0,J_{0},\dotsc,J_{0})\in^{(3n+1)\times(3n+1)}\ \ \text{with}\ \ J_{0}=\left(\begin{smallmatrix}0&-1&0\\ 1&0&-1\\ 0&1&0\end{smallmatrix}\right),$
the dissipation matrix
$R=\text{diag}(R_{0},\dotsc,R_{0},\tfrac{1}{R_{L}})\in^{(3n+1)\times(3n+1)}\ \ \text{with}\ \ R_{0}=\left(\begin{smallmatrix}\tfrac{1}{R_{1}}+\tfrac{1}{R_{d}}&-\tfrac{1}{R_{d}}&0\\ -\tfrac{1}{R_{d}}&\tfrac{1}{R_{d}}&0\\ 0&0&0\end{smallmatrix}\right)$
and the input matrix corresponding to a voltage input $u=\imath$
$B=\begin{pmatrix}1&0&\dotsc&0\end{pmatrix}^{\top}.$
We inspect the time interval [0, $[0,\,5\cdot 10^{-6}]$, and consider the parameters $C=5\cdot 10^{-7},\;C_{R}=5\cdot 10^{-9},\;L=10^{-6},\;R_{d}=5,\;R_{1}=200,\;R_{L}=10,$ and
$\imath(t)=\sin(5\cdot 10^{6}\pi t),\;x(0)=10^{-7}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{pmatrix}1&\ldots&1\end{pmatrix}^{\top}}.$
Furthermore, we choose $n=5$ leading to a 16-dimensional system, however we stress that the results are very robust with respect to adding more elements to the circuit.
We depict the results in Figure 6. In the left column, we utilize 20 refinement cycles and observe that the goal-oriented scheme outperforms the uniform refinements in view of the energy balance. However, in view of the trajectories over time, there is clearly a mismatch in the first coordinate. This mismatch is greatly reduced when performing more refinement cycles as depicted on the right. The reference trajectory is computed with a high-order BDF scheme on a very fine grid using 10001 time steps.
Figure 6. Computational results for the electrical circuit. Left: 20 refinement loops; Right: 60 refinement loops. The vertical axis of the top figures is normalized in view of the total energy supplied to the system.
@@ 7. Conclusion and outlook
We proposed an adaptive and goal-oriented grid refinement approach tailored to the energy balance in port-Hamiltonian systems. To this end, we formulated the state equation and the corresponding adjoint equation as variational problems and proved a result for higher regularity of the adjoint for right-hand sides occurring in the proposed scheme. To guarantee an efficient evaluation of the error estimator, we further provided a block-diagonal and stability-exploiting approximation for the adjoint system which we motivated from a linear algebraic perspective via a block-Jacobi method. Then, we provided a numerical example highlighting that the proposed scheme leads to smaller violations of the energy balanced when compared to step size controlled or higher-order time stepping schemes.
In the future, we plan to consider goal-oriented multirate schemes [9] for pHs and to relate the dissipativity-based arguments of Section 4 with the approach using generalized Green’s functions to capture error transport as proposed in [10]. Furthermore, a combination with a symplectic integrator is expected to lead to a symbiosis of grid adaptivity and geometric integration.